OpenStudy (anonymous):

how do you put into formula 3 times a number + 2 times the second number = 4 times the third number and the have to be integers in a row. I found 14, 15, 16 but how do you show the formula

OpenStudy (anonymous):

and the total of the first two are 8 less than the thrid

OpenStudy (anonymous):

8 more sorry

OpenStudy (anonymous):

put the first number at x, second is x + 1, third is x+2 then \[3x+2(x+1)=4(x+2)\]

OpenStudy (anonymous):

oh wait, total of first two are 8 less than the third would be \[x+x+8=x+2\]

OpenStudy (anonymous):

Let me say it again 3 times the first integer plus two times the second integer is more than 8 four times the third integer

OpenStudy (turingtest):

3 times the first plus 2 times the second is 8 more than 4 times the third?

OpenStudy (turingtest):

if so then\[3x+2(x+1)=4(x+2)+8\]is the formula